# How Does the Textual Information Affect the Retrieval of Multimodal In-Context Learning?

### v1

@@ Abstract
The increase in parameter size of multimodal large language models (MLLMs) introduces significant capabilities, particularly in-context learning, where MLLMs enhance task performance without updating pre-trained parameters. This effectiveness, however, hinges on the appropriate selection of in-context examples, a process that is currently biased towards visual data, overlooking textual information. Furthermore, the area of supervised retrievers for MLLMs, crucial for optimal in-context example selection, continues to be uninvestigated. Our study offers an in-depth evaluation of the impact of textual information on the unsupervised selection of in-context examples in multimodal contexts, uncovering a notable sensitivity of retriever performance to the employed modalities. Responding to this, we introduce a novel supervised MLLM-retriever MSIER that employs a neural network to select examples that enhance multimodal in-context learning efficiency. This approach is validated through extensive testing across three distinct tasks, demonstrating the method’s effectiveness. Additionally, we investigate the influence of modalities on our supervised retrieval method’s training and pinpoint factors contributing to our model’s success. This exploration paves the way for future advancements, highlighting the potential for refined in-context learning in MLLMs through the strategic use of multimodal data.
@@ 1 Introduction
The capability of large language models (LLMs) for in-context learning (ICL) has garnered significant attention and exhibited remarkable efficacy across a diverse range of downstream tasks An et al. (2023); Min et al. (2022b); Wang et al. (2023). ICL aims to adapt the pre-trained model to achieve high performance on downstream tasks without any update of parameters that require extra computing resources for the further training process. To be specific, the primary advantage of in-context learning (ICL) lies in its capacity to assimilate and adapt to new information from its context. This is achieved by utilizing a series of question-and-answer pairs from the training dataset, known as in-context examples, which enable learning without necessitating updates to the model’s parameters. Significantly, recent developments in multimodal large language models (MLLMs), which are constructed based on the foundations of large language models (LLMs), have similarly exhibited capabilities in in-context learning (ICL)Alayrac et al. (2022); Peng et al. (2023); Zhao et al. (2023); Yin et al. (2023), termed as multimodal in-context learning (M-ICL).
Figure 1: An overview of multimodal in-context example retrieval: This process involves receiving an image or an image-text query from the test dataset, and then using a retrieval mechanism to find similar examples in a training dataset. These examples and the original query (collectively called the prompt) are then inputted into a multimodal large language model (MLLM) to generate the output.
@@
In contrast to Natural Language Processing (NLP) Dong et al. (2023); Coda-Forno et al. (2023), where the primary input is text, multimodal tasks predominantly utilize a single image (image-text pair) as input and the memory for retrieval consists of multimodal examples. Current methodologies for unsupervised retrieval of in-context examples in multimodal large language models (MLLMs) predominantly focus on visual informationAlayrac et al. (2022); Yang et al. (2022); Gu et al. (2023), thereby overlooking the significance of linguistic data. This oversight prompts inquiries regarding the impact of textual content on retrieval processes and whether the inclusion of additional modalities could potentially enhance the efficacy of MLLMs’ in-context learning capabilities. Moreover, in the pursuit of developing a supervised retrieval framework, the question arises as to whether the integration of textual information constitutes a crucial element. Despite the importance of these considerations, they have received minimal attention within the realm of M-ICL so far.
In the work, our attention is centered on multimodal in-context learning, an emerging concept in the field with limited existing research on its practical application. For the first time, we undertake a thorough investigation into the effects of textual information for retrieval of in-context examples on MLLM. Through this research, we have identified a pivotal issue: the performance of the selection of in-context examples on downstream tasks is notably influenced by the selection of modality for both unsupervised and supervised retrievers. The contributions of this paper are summarized as follows:
  • We present the first analysis into the role of textual modal in the unsupervised retrieval of in-context examples for M-ICL and demonstrate that the addition of textual modal plays a crucial role in the improvement of M-ICL performance.
  • By taking both images and texts into consideration when selecting in-context examples, we further design a Multimodal Supervised In-context Examples Retrieval (MSIER) framework with more efficient example selection performance and explore the impact of text modal in the training of the proposed supervised retriever.
  • Extensive experiments on three typical multimodal tasks (image captioning, visual question answering, and rank classification) demonstrate the high efficiency of our constructed unsupervised and supervised retrieval methods that achieve the best performance. Besides, we also provide valuable insights into the importance of various modalities in selecting good in-context examples for M-ICL.
@@ 2 Related Work
In-context Learning In-context learning was first proposed by Brown et al. (2020) in their paper introducing GPT-3. It is a significant departure from the traditional learning method based on stochastic gradient descent and does not involve any parameter updates: the model needs to be provided with a few examples of the task as part of the context for text generation. The size of the model and training data were thought to be key to training a model with in-context learning capabilities. More recently, there has been more research on the exact causes of in-context learning. Min et al. (2022a) has proposed MetaICL, a meta-training framework to elicit in-context learning capabilities in text-only language models. MetaICL conditions each example with related in-context examples during training. In the context of MLLM, ICL has been extended to more modalities, leading to Multimodal ICL (M-ICL) Alayrac et al. (2022); Li et al. (2023); Chen et al. (2023b).
@@
In-context Examples Retrieval in NLP and CV The field of natural language processing has identified that the selection of in-context examples significantly influences performance, as evidenced by Agrawal et al. (2022) and Min et al. (2022b). Furthermore, the construction of these in-context examples, often referred to as prompts, including aspects such as relevance and diversity of retrieved examples, has been reported to impact performance as well. This understanding has guided the community toward investigating optimal methods for selecting effective in-context examples for large language models, an area that has also stimulated our research. In their study, Liu et al. (2021) posited that effective in-context examples should bear semantic similarity to query sentences. Based on this hypothesis, they advocated for the selection of the nearest neighbors in the training set, as determined by a sentence encoder like RoBERTa Liu et al. (2019). Rubin et al. (2022) initially employed an unsupervised approach to gathering potential candidates, from which the most suitable examples were selected using a supervised method.
@@
Correspondingly, Zhang et al. (2023) explored the application of a supervised retriever within the scope of visual in-context learning, demonstrating enhanced performance outcomes. Nevertheless, the significance of textual information is overlooked within the framework of unsupervised retrieval for multimodal in-context learning, and the impact of various modalities on the development of the supervised retriever continues to be an area requiring investigation.
@@
Retrieval in Multimodal Models Retrieval augmentation for multimodal models enhances their capabilities by infusing externally retrieved information into their workflow. Chen et al. (2022) proposed MuRAG, which accesses an external multimodal memory to augment language generation. Yasunaga et al. (2023) presented the first retrieval-augmented multimodal model that enables a base model to refer to relevant text and images fetched from external memory. Ramos et al. (2023) introduced a new approach to image captioning that generates sentences given the input image and retrieved captions. However, these methods require further training or fine-tuning to synthesize aspects from both the input and retrieved information. On the contrary, the in-context examples provided for M-ICL are used as part of the input sequence to the model, without deeper integration into the model’s architecture that requires further parameter updates, allowing the model to interpret and generate based on its existing capabilities.
@@ 3 Method
Figure 2: Overview of the Multimodal Supervised In-context Examples Retrieval (MSIER) Method: The fundamental principle involves assessing the in-context learning performance for each source instance, thereafter identifying and choosing those instances exhibiting the most favorable or least favorable outcomes. These selected instances are then utilized to form a dataset, categorized as either positive or negative, which is essential for the facilitation of contrastive learning.
@@ 3.1 Multimodal In-Context Learning
Within the realm of multimodal tasks such as image captioning, a pre-trained Multimodal Language and Learning Model (MLLM) utilizes an image and its corresponding caption as an in-context example to delineate requirements of the task — essentially instructing the model on the expected form of input and output. Subsequently, when presented with a new image, the model can produce a more precise caption based on these learned patterns. Crucially, Multimodal In-Context Learning (M-ICL) maintains parameters of the pre-trained model unchanged, thereby offering a more resource-efficient approach for tackling downstream tasks.
@@ 3.2 Importance of text information for unsupervised MLLM retrieval
As for the framework of the Multimodal Unsupervised In-context Examples Retrieval (MUIER) which is a prevalent retrieval methodology, our analysis initially focuses on the contribution of various modalities. Traditionally, existing MLLMs that present outstanding M-ICL capability have primarily utilized image data for the facilitation of M-ICL tasks. For instance, Flamingo Alayrac et al. (2022) employed the Retrieval-based In-Context Example Selection (RICES) strategy Yang et al. (2022) for the identification of appropriate in-context examples, predominantly assessing the similarity between the query image and the images stored in the memory (training dataset). Nonetheless, this approach overlooks the significant role of textual information, which indeed influences the efficacy of M-ICL.
To validate the effectiveness of text-augmented MUIER, we performed a comprehensive comparison across diverse configurations of unsupervised retrievers, encompassing three main settings specifically designed for the image captioning task: (1) Q-I-M-I (Query-Image-Memory-Image) indicates the case where only image information of image-text pairs preserved in the memory was applied for the retrieval of context. As described in RICES, we only consider the cosine similarity between the query image $q_{i}$ and the memory image $m_{i}$ based on vision features extracted by the unsupervised retriever. (2) Q-I-M-IT describes the standard setting for MUIER by pairing the memory image with its corresponding label. (3) Q-I-M-IT w/ mask pertains to the scenario in which a subset of words (N=3) within the label text are removed, thereby rendering only a fraction of the words within the label text accessible to the MLLM during the evaluation phase (to determine if a sentence fragment is sufficient for extracting textual features).
@@
Figure 3 presents the M-ICL performance in different unsupervised retriever settings, given the same query (Q) and memory (M) configuration. Compared with the standard setting, both the Q-I-M-IT w/ mask and Q-I-M-IT settings exert a positive impact on the M-ICL performance to varying degrees, which verifies the substantial influence of textual content on M-ICL performance.
@@ 3.3 Multimodal Supervised Prompt Retriever
The unsupervised methodology depicted earlier is not explicitly designed for multimodal in-context learning applications. Instead, its efficacy relies on the preliminary training phase of the feature extraction mechanism, where the objective function utilized may not align with the demands of multimodal in-context learning scenarios. In contrast, we propose a distinct strategy that is based on supervised in-context examples retrieval, assuming the availability of labeled training data as shown in Figure 2. The primary aim of this approach is to refine the original retriever, denoted as $R$, to ensure that the chosen in-context example(s) contribute effectively towards enhancing the log-likelihood maximization.
It is important to acknowledge that evaluating each candidate within the dataset for MLLM training entails a significant expenditure of time and computational resources. This is attributed to the extensive size of existing MLLMs and the substantial overhead associated with preprocessing image data. Consequently, in alignment with the methodologies outlined in Rubin et al. (2022), we initially narrowed down the Top-$K$ ($K$=50) candidates by assessing the cosine similarity of multimodal information through the application of the unsupervised retriever. Drawing upon the findings presented in 3.2, our approach integrates both visual and textual data for the identification of Top-$K$ examples, which are subsequently employed in the training of our proposed multimodal in-context examples retriever.
@@
Training Adopting the contrastive learning framework as mentioned by Rubin et al. (2022), we introduce a query encoder designed to process inputs comprising image or image-question pairs, and a context encoder tasked with handling candidate prompts represented as image-prompt pairs, for feature extraction aimed at subsequent optimization. Both encoders are initialized with vision encoders and text encoders of CLIP model Radford et al. (2021). During each iteration, a minibatch $B$ is selected from the training dataset to fine-tune the CLIP model. Furthermore, for every instance within the assembled memory $B$ with $N$ image-text pair, both a positive and a negative example are independently sampled from the Top-$K$ positive and negative candidates. This sampling strategy is implemented to develop a similarity metric. Specifically, this metric should ensure that for a given test example $x_{q}$, it exhibits similarity to training instances that facilitate the decoding of $y_{q}$. The formulation of the contrastive loss is executed as follows:
@@
3.4 Investigating the Importance of Textual Information in Supervised Retriever
In Section 3.2, the significance of various modalities is examined within the context of unsupervised retrieval, yet the impact of textual data on MSIER has not been investigated. To thoroughly assess the contribution of text in the training of supervised retrievers, we conduct a detailed comparison of two scenarios involving fine-tuned CLIP models, distinguishing them based on their approach to incorporating textual information within the image captioning task.
We decouple the training and evaluation settings to investigate the significance of textual data in our proposed supervised retrieval framework. Specifically, we employ distinct query-memory configurations during the retriever’s training phase and evaluation phase. As depicted in Figure 4, two training configurations – Q-I-M-I and Q-I-M-IT – were implemented, and their performance was assessed under two corresponding settings to ensure an equitable comparison. Within the context of the Q-I-M-I setting, textual data during the training phase has a negligible influence on the retriever’s performance. In contrast, incorporating textual information during the training of the supervised multimodal retriever markedly enhances its effectiveness, resulting in a significant performance improvement.

### v2

@@ Abstract
The increase in parameter size of multimodal large language models (MLLMs) introduces significant capabilities, particularly multimodal in-context learning, where MLLMs enhance task performance without updating pre-trained parameters. However, this effectiveness hinges on the appropriate selection of in-context examples, a process currently biased towards visual data, overlooking textual information. More importantly, the area of supervised retrievers for retrieval of multimodal in-context learning, crucial for optimal in-context example selection, continues to be uninvestigated. Our study provides an in-depth evaluation of the impact of textual information on the unsupervised selection of in-context examples in multimodal contexts, uncovering a notable sensitivity of retriever performance to the employed modalities. Based on the above finding, we introduce a novel supervised MLLM prompt retriever MSIER that leverages a trained retriever based on MLLM’s confidence to select examples, which enhances multimodal in-context learning efficiency. This approach is validated through extensive testing across three different tasks, demonstrating the method’s effectiveness. Additionally, we investigate the influence of modalities on our supervised retrieval method’s training and explore the transferability of the supervised prompt retriever. This exploration paves the way for future advancements, highlighting the potential for refined in-context learning in MLLMs through the strategic use of multimodal data. The public code is available at https://github.com/NUS-HPC-AI-Lab/Multimodal-ICL-Retriever.
@@ 1 Introduction
The capability of large language models (LLMs) for in-context learning (ICL) has acquired significant attention and exhibited remarkable efficacy across a diverse range of downstream tasks An et al. (2023); Min et al. (2022b); Wang et al. (2023). ICL aims to adapt the pre-trained model to achieve high performance on downstream tasks without any update of parameters that require extra computing resources for the further training process. To be specific, the primary advantage of in-context learning lies in its capacity to assimilate and adapt to new information from its context. This is achieved by utilizing a series of question-and-answer pairs from the training dataset, known as in-context examples, which enable learning without necessitating updates to the model’s parameters. Significantly, recent developments in multimodal large language models (MLLMs), which are constructed based on the foundations of large language models, have similarly exhibited capabilities in in-context learning Alayrac et al. (2022); Peng et al. (2023); Zhao et al. (2023); Yin et al. (2023), termed as multimodal in-context learning (M-ICL).
Figure 1: An overview of multimodal in-context example retrieval: This process involves receiving an image or an image-text query from the test dataset, and then using a retrieval mechanism to find similar examples in a training dataset. These examples and the original query (collectively called the prompt) are then inputted into a MLLM to generate the output.
@@
In contrast to Natural Language Processing (NLP) Dong et al. (2023); Coda-Forno et al. (2023), where the primary input is text, multimodal tasks predominantly utilize a single image (image-text pair) as input and the memory for retrieval consists of multimodal examples. Current methodologies for unsupervised retrieval of in-context examples in multimodal large language models (MLLMs) predominantly focus on visual informationAlayrac et al. (2022); Yang et al. (2022); Gu et al. (2023), thereby overlooking the significance of linguistic data. This oversight prompts inquiries regarding the impact of textual content on retrieval processes and whether the inclusion of additional modalities could potentially enhance the efficacy of MLLMs’ in-context learning capabilities. Moreover, the supervised retrieval framework for M-ICL has received minimal attention so far.
In this work, our attention is centered on M-ICL, an emerging concept in the MLLMs field with limited existing research on its practical application. Specifically, we undertook a thorough investigation into the effects of textual information for retrieval of in-context examples in M-ICL. Through this research, we have developed a supervised retriever that aligns with existing knowledge from MLLMs. This framework achieves high efficiency and demonstrates excellent transferability across various datasets and sizes of MLLMs. The contributions of this paper are summarized as follows:
  • We present the first thorough analysis into the role of textual modality in both unsupervised and supervised retrieval of in-context examples for M-ICL and demonstrate that the addition of textual modality plays a crucial role in improving M-ICL performance.
  • By considering both visual and textual information when selecting in-context examples, we design a Multimodal Supervised In-context Examples Retrieval (MSIER) framework with more efficient example selection performance by using a foundation MLLM scorer to evaluate the relevance and suitability of potential examples from multimodal data.
  • Extensive experiments on three typical multimodal tasks demonstrate the high efficiency of our constructed supervised retrieval method that achieves the best performance. Besides, we further provide valuable insights regarding the importance and transferability of the supervised method in selecting highly relevant in-context examples for M-ICL.
@@ 2 Related Work
In-context Examples Retrieval in NLP and CV The field of natural language processing has identified that the choice of in-context examples significantly influences performance, as evidenced by Agrawal et al. (2022) and Min et al. (2022b). Furthermore, the construction of these in-context examples, often referred to as prompts, including aspects such as relevance and diversity of retrieved examples, has been reported to impact performance as well. This understanding has guided the community toward investigating optimal methods for selecting effective in-context examples for large language models. In their study, Liu et al. (2021) posited that effective in-context examples should bear semantic similarity to query sentences. Based on this hypothesis, they advocated for the selection of the nearest neighbors in the training set, as determined by a sentence encoder like RoBERTa Liu et al. (2019). Rubin et al. (2022) initially employed an unsupervised approach to gathering potential candidates, from which the most suitable examples were selected using a supervised method.
@@
Correspondingly, Zhang et al. (2023) explored the application of a supervised retriever within the scope of visual in-context learning, demonstrating enhanced performance outcomes. Nevertheless, the significance of textual information is overlooked within unsupervised and supervised retrievals for M-ICL, and the impact of various modalities on the development of the supervised retriever continues to be an area requiring investigation.
@@
Multimodal Retrieval-augmented Generation Retrieval augmentation for multimodal models enhances their capabilities by infusing externally retrieved information into their workflow. Chen et al. (2022) proposed MuRAG, which accesses an external multimodal memory to augment language generation. Yasunaga et al. (2023) presented the first retrieval-augmented multimodal model that enables a base model to refer to relevant text and images fetched from external memory. Ramos et al. (2023) introduced a new approach to image captioning that generates sentences given the input image and retrieved captions. The retrieval process in RAG is designed to retrieve supplementary information to enhance the context for text generation and to produce more informed and contextually rich outputs. On the other hand, the retrieval of in-context examples in M-ICL aims to find examples that demonstrate how to perform a given task, providing explicit guidance to the model on how to generate its output. However, the role of textual information in the task of in-context example retrieval has remained unexplored so far.
In-context Learning In-context learning was first proposed by Brown et al. (2020) in their paper introducing GPT-3. It is a significant departure from the traditional learning method based on stochastic gradient descent and does not involve any parameter updates: the model needs to be provided with a few examples of the task as part of the context for text generation. The size of the model and training data were thought to be key to training a model with in-context learning capabilities. More recently, there has been more research on the exact causes of in-context learning. Min et al. (2022a) has proposed MetaICL, a meta-training framework to elicit in-context learning capabilities in text-only language models. MetaICL conditions each example with related in-context examples during training. In the context of MLLMs, ICL has been extended to more modalities, leading to Multimodal ICLAlayrac et al. (2022); Li et al. (2023); Chen et al. (2023b).
@@ 3 Method
Figure 2: Overview of the MSIER Method: The fundamental principle involves assessing the in-context learning performance for each source instance, thereafter identifying and choosing those instances exhibiting the most favorable or least favorable outcomes. These selected instances are then utilized to form a dataset, categorized as either positive or negative, which is essential for the facilitation of contrastive learning. The examples with high CIDEr scores (corresponding to low NLL loss during the scoring process) are selected as positive samples.
@@ 3.1 Multimodal In-Context Learning
Within the realm of multimodal tasks such as image captioning, a pre-trained MLLM utilizes an image and its corresponding caption as an in-context example to delineate requirements of the task — essentially instructing the model on the expected form of input and output. Subsequently, when presented with a new image, the model can produce a more precise caption based on these learned patterns. Crucially, M-ICL maintains the parameters of the pre-trained model unchanged, thereby offering a more resource-efficient approach for tackling downstream tasks.
@@ 3.2 Importance of text information for unsupervised MLLM retrieval
As for the framework of the Multimodal Unsupervised In-context Examples Retrieval (MUIER) which is a prevalent retrieval methodology, our analysis initially focuses on the contribution of various modalities. Traditionally, existing MLLMs that present outstanding M-ICL capability have primarily utilized image data to facilitate M-ICL tasks. For instance, Flamingo Alayrac et al. (2022) employed the Retrieval-based In-Context Example Selection (RICES) strategy Yang et al. (2022) for the identification of appropriate in-context examples, predominantly assessing the similarity between the query image and the images stored in the memory (training dataset). Nonetheless, this approach overlooks the significant role of textual information, which indeed influences the efficacy of M-ICL.
To validate the effectiveness of text-augmented MUIER, we performed a comprehensive comparison across diverse configurations of unsupervised retrievers, encompassing three main settings specifically designed for the image captioning task: (1) Q-I-M-I (Query-Image-Memory-Image) indicates the case where only image information of image-text pairs preserved in the memory was applied for the retrieval of context. As described in RICES, we only consider the cosine similarity between the query image and the memory image based on vision features extracted by the unsupervised retriever. (2) Q-I-M-IT describes the standard setting for MUIER by pairing the memory image with its corresponding text label. Specifically, when we use an image as a query, we will calculate its similarity between the query image $i_{q}$ and the memory image $i_{m}$ as well as its caption $c_{m}$, which means the final similarity $S=cos(i_{q},i_{m})+cos(i_{q},c_{m})$.
@@
Figure 3 presents the M-ICL performance in different unsupervised retriever settings, given the same query (Q) and memory (M) configuration. Compared with the standard setting, the Q-I-M-IT setting exerts a positive impact on the M-ICL performance to varying degrees, which verifies the substantial influence of textual content on M-ICL performance.
@@ 3.3 Multimodal Supervised Prompt Retriever
The unsupervised methodology depicted earlier is not explicitly designed for multimodal in-context learning applications. Instead, its efficacy relies on the preliminary training phase of the feature extraction mechanism, where the objective function utilized may not align with the demands of multimodal in-context learning scenarios. In contrast, we propose a novel strategy that is based on supervised in-context examples retrieval, assuming the availability of labeled training data as shown in Figure 2. The primary aim of this approach is to refine the original retriever to ensure that the chosen in-context example(s) contribute effectively towards enhancing the log-likelihood maximization.
It is important to acknowledge that evaluating each candidate within the dataset for MLLM training entails a significant expenditure of time and computational resources. This is attributed to the extensive size of existing MLLMs and the substantial overhead associated with preprocessing image data. Consequently, in alignment with the methodologies outlined in Rubin et al. (2022), we initially narrowed down the Top-$N$ ($N$=50) candidates by assessing the cosine similarity of multimodal information through the application of the unsupervised retriever. Drawing upon the findings presented in 3.2, our approach integrates both visual and textual data for the identification of Top-$N$ examples, which are subsequently employed in the training of our proposed multimodal in-context examples retriever.
@@
Training Adopting the contrastive learning framework as mentioned by Rubin et al. (2022), we introduced a query encoder designed to process inputs comprising image or image-question pairs, and a context encoder tasked with handling candidate prompts represented as image-prompt pairs, for feature extraction aimed at subsequent optimization. Both encoders are initialized with vision encoders and text encoders of CLIP model Radford et al. (2021). During each iteration, a minibatch $B$ is selected from the training dataset to fine-tune the CLIP model. Furthermore, for every instance within the assembled memory $B$ with $N$ image-text pair, both a positive and a negative example are independently sampled from the Top-$K$ positive and negative candidates. This sampling strategy is implemented to develop a similarity metric. Specifically, this metric should ensure that for a given test example $x_{q}$, it exhibits similarity to training instances that facilitate the decoding of $y_{q}$. The formulation of the contrastive loss is executed as follows:
@@
3.4 Importance of Textual Information in Supervised Retriever
In Section 3.2, the significance of various modalities is examined within the context of unsupervised retrieval, yet the impact of textual data on MSIER has not been investigated. To thoroughly assess the contribution of text in the training of supervised retrievers, we conducted a detailed comparison of two scenarios involving fine-tuned CLIP models, distinguishing them based on their approach to incorporating textual information within the image captioning task.
We decoupled the training and evaluation settings to investigate the significance of textual data in our proposed supervised retrieval framework. Specifically, we employed distinct query-memory configurations during the retriever’s training phase and evaluation phase. As depicted in Figure 4, two training configurations – Q-I-M-I and Q-I-M-IT – were implemented, and their performance was assessed under two corresponding settings to ensure an equitable comparison. Within the context of the Q-I-M-I setting, textual data during the training phase has a negligible influence on the retriever’s performance. In contrast, incorporating textual information during the training of the supervised multimodal retriever markedly enhances its effectiveness, resulting in a significant performance improvement.